OpenStudy (loser66):

Find all solutions of the equation \(x^2 +x-6=0\) in the ring \(Z_{14}\) by factoring the quadratic polynomial. Please, help

OpenStudy (anonymous):

\(x^2+x-6=x^2+3x-2x-6=x(x+3)-2(x+3)=(x-2)(x+3)\)

OpenStudy (anonymous):

note \((x-2)(x+3)\equiv0\pmod{14}\) but \(14=1\cdot14=2\cdot7\) ergo we have multiple cases: 1. we have \(x+3\equiv0\pmod{14}\) 2. we have \(x-2\equiv 0\pmod{14}\) 3. we have \(x-2\equiv0\pmod2\) and \(x+3\equiv0\pmod7\) 4. we have \(x-2\equiv0\pmod7\) and \(x+3\equiv0\pmod2\)

OpenStudy (anonymous):

notice that since \(14=2\times7\) is not prime we have two zero divisors \(2,7\) and therefore \(\mathbb{Z}/14\mathbb{Z}\) is not an integral domain -- this explains why the bottom two possibilities must now be considered

OpenStudy (anonymous):

in familiar rings like \(\mathbb{Z},\mathbb{Q},\mathbb{R},\mathbb{C}\) we have no zero divisors i.e. \(ab=0\implies (a=0)\lor (b=0)\) -- these are known as integral domains. our case here does indeed have zero divisors so \(ab=0\) actually gives MORE such possibilities

OpenStudy (anonymous):

we know that in \(\mathbb{Z}/14\mathbb{Z}\) that \(a=0\implies ab=0\) but we also know that \(a=2,b=7\implies ab=0\); the existence of zero divisors means we have more possibilities

OpenStudy (loser66):

so, we just solve x =-3 (mod 14) x = 2 (mod 14) right?

OpenStudy (anonymous):

that only covers the possibilities \((1),(2)\) -- there are more possibilities than these... P.S. \(x\equiv-3\equiv11\pmod{14}\) which is the element in \(\mathbb{Z}/14\mathbb{Z}\)

OpenStudy (anonymous):

you must also solve \((3),(4)\)

OpenStudy (loser66):

Oh !! I understand the whole thing now. Since this problem relate to the previous one, which stated that although (x-a) (x+a) =0 , we cannot conclude that (x-a) =0 or (x+a) =0 i understood that concept. But to this problem, I don't know how to solve with "in Z-14"

OpenStudy (loser66):

so, after solving 1,2,3,4 combine the solutions are the final answer, right?

OpenStudy (anonymous):

dunno what you mean by combine... each of 1,2,3,4 gives distinct solutions

OpenStudy (anonymous):

take the union of their solution sets sure

OpenStudy (loser66):

got you, union, that my "combine", Thank you very much.

OpenStudy (anonymous):

1. we have \(x+3\equiv0\pmod{14}\) as we noted, this gives \(x\equiv-3\equiv 11\pmod{14}\) so \(x=11\) is our solution in \(\mathbb{Z}/14\mathbb{Z}\) 2. we have \(x−2\equiv0\pmod{14}\) as we noted, this gives \(x\equiv2\pmod{14}\) so \(x=2\) is our solution in \(\mathbb{Z}/14\mathbb{Z}\) 3. we have \(x−2\equiv0\pmod2\) and \(x+3\equiv0\pmod7\) notice \(x-2\equiv x\equiv 0\pmod2\) ergo \(x\in\{0,2,4,6,8,10,12\}\subset\mathbb{Z}/14\mathbb{Z}\) also we have \(x\equiv-3\equiv4\pmod7\) ergo \(x\in\{4,11\}\subset\mathbb{Z}/\mathbb{14Z}\) the solutions that satisfy both are then \(\{0,2,4,6,8,10,12\}\cap\{4,11\}=\{4\}\) 4. we have \(x−2\equiv0\pmod7\) and \(x+3\equiv0\pmod2\) notice \(x\equiv2\pmod7\) ergo \(x\in\{0,2,9\}\subset\mathbb{Z}/14\mathbb{Z}\) also we have \(x\equiv-3\equiv1\pmod2\) ergo \(x\in\{1,3,5,7,9,11,13\}\subset\mathbb{Z}/\mathbb{14Z}\) the solutions that satisfy both are then \(\{0,2,9\}\cap\{1,3,5,7,9,11,13\}=\{9\}\)

OpenStudy (anonymous):

therefore the total set of solutions is the union of all of the above solutions \(\{11\}\cup\{2\}\cup\{4\}\cup\{9\}=\{2,4,9,11\}\)

OpenStudy (loser66):

Woaahahha . thanks a ton.

OpenStudy (anonymous):

this idea is the heart of the Chinese remainder theorem: http://en.wikipedia.org/wiki/Chinese_remainder_theorem#Statement_for_general_rings